[Hallc_running] [New Logentry] RC daily update

simona at jlab.org simona at jlab.org
Wed Jul 17 11:20:08 EDT 2019

Logentry Text:
We got 5 pass beam in Hall C during owl around 3:30 am. Harp scans and beam positions looked good: https://logbooks.jlab.org/entry/3705344 , https://logbooks.jlab.org/entry/3705345 , https://logbooks.jlab.org/entry/3705362

We started production soon after at ~13 muA (that's the beam current that the Hall needed/requested for now). At the 7:45 meeting I was told that there will be an interruption in Hall C beam delivery today to tune to B. The expectation is that we can resume production running in C by 1 pm. 

We expect to finish data taking at the current Q2 point sometime during owl. If that's the case, next we will request a beam energy measurement and a BCM calibration. We will make an access tomorrow during the day shift to rotate HMS (and SHMS).

There will be an RC meeting today at 4:30 pm on the second floor of the counting house: https://bluejeans.com/861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3705515
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190717/4cc690f7/attachment.html>

More information about the Hallc_running mailing list